OpenStudy (charlotte123):

Brian is creating a collage on a piece of cardboard that has an area of 150r^2 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of sqrt r^3 square centimeters. Use the given information to determine an expression for the total number of pieces of paper used.

hartnn (hartnn):

total area of cardboard/ area of each piece = total number of pieces sounds logical ?

OpenStudy (charlotte123):

@hartnn yeah :D so for the expression, should it be 150r^2/√r^3.

OpenStudy (charlotte123):

Im not sure D: its a pretest so when it comes to things I havent studied yet, I suck T_T

hartnn (hartnn):

yes, thats correct :) but you need to simplify that.

OpenStudy (charlotte123):

hmmm....so I cant put that for the equation? :O

hartnn (hartnn):

no...

OpenStudy (charlotte123):

@hartnn ahh so it needs to simplified ^-^ would it be 1501√r? :) sorry if Im wrong, still learning this type of thing >.>

OpenStudy (charlotte123):

oops the 1 got in their >.> I meant 150√r :)

OpenStudy (anonymous):

You need to divide the total area by the area of each sheet of paper.

hartnn (hartnn):

\(\sqrt{r^3}=\sqrt{r*r*r}=r\sqrt r\) so, \(150/\sqrt r\) is correct :)

OpenStudy (charlotte123):

@love but for the problem, I need an equation for the information that was given >.>

OpenStudy (anonymous):

150(r^2)/(r^3)^0.5

OpenStudy (charlotte123):

@hartnn :O so the equation should be 150√r? :D

OpenStudy (charlotte123):

@hartnn lol I mean 150 divided by √r :)

OpenStudy (charlotte123):

@love where did the 0.5 come from? :O

hartnn (hartnn):

yes, yes, 150√r not what i wrote :P

OpenStudy (charlotte123):

@hartnn so 150√r :D lol the equation is shorter than I thought it was :O thanks for ur help! :D

hartnn (hartnn):

welcome ^_^

OpenStudy (charlotte123):

@hartnn and thanks for the medal :D ttyl ~(^O^)~